# Expanding Horizons in Short Text Analysis: Integrating LLMs and VAEs for Enhanced Topic Modeling

### v2

@@ 4.3 Implementation Details
For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper. Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model. This checkpoint, open-sourced by the original authors, has been warmed up and fine-tuned on multiple datasets to function as the query rewriter. Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine. These trainings were conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
@@
Retrieval  For our web search engine, we opt for the Brave Search Engine, which, although it may provide slightly lower quality results compared to major competitors like Google or Bing, offers a significantly more cost-effective API. This search API retrieves website snippets, simulating a typical user experience of entering a query in a search engine, pressing Enter, and reviewing the top results at a glance. For local retrieval, we utilize WikiDPR, a specialized subset of Wikipedia collections tailored for the Dense Passage Retrieval (DPR) model DPR. This database consists of 21 million passages from Dec. 20, 2018, each limited to 100 words, along with their 768-dimensional embedded vectors. The retrieval process involves converting a query into a DPR embedding and finding the top k vectors with the closest L2 distances. For both systems, we retrieve the top 5 results, concatenate them with the original query, and feed them to the LLM reader.
Generation  Although different prompting strategies may influence the performance of the question-answering task, this aspect is not the primary focus of our study, so we adhere to the same answer prompts used in the RRR rrr framework. The prompts we used are detailed in Table 1.
Trainable Scheme  For Trainable ERRR, we employ T5-Large t5, an open-source model with 770 million parameters, as the query optimizer. We fine-tune this student model using knowledge distillation from GPT-3.5-Turbo. The distillation dataset was assembled by selecting questions from training sets of each QA dataset, with GPT-3.5-Turbo generating the responses under identical settings utilized in the frozen scheme. We also devised a short eliciting prompt, "Rewrite better search queries to acquire or validate the knowledge needed for the question:", serving as an instruction prefix to guide T5 to adapt to the task. To ensure optimal task-specific outcomes, separate T5 models were trained with 3 epochs for each QA dataset, with a learning rate of 1e-4 and a batch size of 4.
[…]
@@ 4.4 Result
The experimental results across three datasets and two different retrieval tools are presented in Table 2 and Table 3. The Frozen ERRR framework outperforms all the baseline methods—Direct, Frozen RRR, and Trainable RRR—regardless of the retrieval system. This underscores the effectiveness of addressing the pre-retrieval gap and demonstrates adaptability across different retrieval systems and data sources. Additionally, to our surprise, the Trainable ERRR framework further outperforms all other baselines and even its teacher model across all three datasets. We attribute this superior performance to the distillation process, which not only allows the student model to learn from the teacher model’s insights but also to generalize more effectively. In this case, the student model compresses the teacher’s knowledge, potentially learning to focus on critical features and patterns while discarding less relevant information. This compression likely leads to such performance, enabling the model to adapt more proficiently to specific optimization tasks.
[…]
@@
Nevertheless, the impact of the ERRR framework is more noticeable on web search retrieval systems, as evidenced by the degree of performance enhancement between web search and dense retrievers detailed in Table 2 and Table 3. Moreover, the results reveal that both Frozen RRR and Trainable RRR underperform the direct method in the PopQA and HotPotQA datasets when using a dense retriever. This underperformance is attributed to the low quality of the results retrieved from the static corpus, which only includes Wikipedia passages with limited length and knowledge span. Additionally, the low quality of retrieval contributes to the retrieval of more irrelevant documents, which subsequently distracts the Large Language Model (LLM) from correctly answering questions. However, such distraction does not cause ERRR to underperform baseline methods. This suggests that the query optimizer mainly optimizes queries to align with the LLM’s informational needs, thereby reducing the likelihood of retrieving too many irrelevant passages. This is particularly beneficial when operating with a document collection of sub-optimal quality. A relevant case study can be found in Appendix A.
@@ 4.5 Cost and Latency
Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks. To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct react and Self-RAG selfrag. This experiment was carried out on 200 randomly selected questions from HotPotQA. The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks. Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby saving on an additional LLM call to GPT-3.5-Turbo.
@@ 5 Conclusion
In this paper, we present Extract-Refine-Retrieve-Read (ERRR) framework for Retrieval-Augmented Generation (RAG) systems. The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness. Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA ambigqa, PopQA popqa and HotPotQA hotpotqa, utilizing both web search tools and a dense retriever with local static corpus. It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems. This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems. Additionally, we have developed and implemented a trainable scheme for the ERRR framework. This approach is both cost-effective and efficient as it relies on only a fine-tuned T5 model trained on a moderately sized dataset and surpasses the performance of the frozen GPT-3.5-Turbo.
@@ 6 Limitation
We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGselfrag and CRAGCRAG. These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands. Due to computational constraints within our study, we focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model. This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) schulman2017proximal, which constrained the potential upper limit of our model’s performance.
@@
{tabularx}
[…]
RRR rewritten query: "actor comedian "The Office" Stories USA cast"
Output: "Ricky Gervais" (incorrect)
Parametric Knowledge Extraction: Stories USA starred actor and comedian Steve Carell from "The Office." Steve Carell is best known for his role as Michael Scott on the hit TV show "The Office," but he has also appeared in a number of films and other television shows. In Stories USA, Carell plays a man who is struggling to come to terms with his past and find his place in the world. The film is a collection of short stories that explore different aspects of the American experience, and Carell’s performance is praised for its depth and emotional resonance. Overall, Stories USA is a thought-provoking and engaging film that showcases Carell’s talent as both a comedic and dramatic actor.
ERRR rewritten queries: "actor and comedian from "The Office" in Stories USA" "Steve Carell role in Stories USA"
Output: "Steven John Carell" (correct)
[…]
RRR rewritten query: "Pakistani actor writer Islamabad Coke Kahani 2012"
Output: "Ali Abbas"(incorrect)
Parametric Knowledge Extraction: One of the Pakistani actors and writers from Islamabad who helped write for the 2012 Pakistani comedy drama sitcom, "Coke Kahani," is Faisal Rehman. Faisal Rehman is a talented actor, writer, and director in the Pakistani entertainment industry. He has appeared in numerous television dramas and films, showcasing his versatility and acting skills. In addition to his acting career, Faisal Rehman has also written scripts for various television shows, including "Coke Kahani," which was well-received by audiences for its humor and engaging storyline. ERRR rewritten queries: "Pakistani actor and writer from Islamabad who helped write for Coke Kahani" "Faisal Rehman contributions to Coke Kahani"
Output: "Yasir Hussain"(correct)
[…]
Table 5: Case Study from RRR and ERRR framework.

### v3

@@ 4.3 Implementation Details
For all baselines, we utilized GPT-3.5-Turbo as the primary LLM and adhered to their implementation from the original paper. GPT-3.5 Turbo was chosen for its balance of performance and cost, aligning with our focus on optimizing retrieval-augmented generation systems rather than benchmarking generative models themselves. While GPT-4 offers improved capabilities, our emphasis remained on augmenting the model’s utility through query optimization. Notably, for the Trainable RRR, we employed the supervised fine-tuned T5 model checkpoint as the base model. This checkpoint, open-sourced by the original authors, has been warmed up and fine-tuned on multiple datasets to function as the query rewriter. Then we replicated their reinforcement learning process since we replaced the original search tool with the Brave Search Engine. These trainings were conducted on the first 1000 data points for each dataset evaluated, with The training parameters set as follows: a learning rate of 2e-5, 3 epochs, and a batch size of 8.
@@
Retrieval  For our web search engine, we opt for the Brave Search Engine, which, although it may provide slightly lower quality results compared to major competitors like Google or Bing, offers a significantly more cost-effective API. This search API retrieves website snippets, simulating a typical user experience of entering a query in a search engine, pressing Enter, and reviewing the top results at a glance. For local retrieval, we utilize WikiDPR, a specialized subset of Wikipedia collections tailored for the Dense Passage Retrieval (DPR) model Karpukhin et al. (2020). This database consists of 21 million passages from Dec. 20, 2018, each limited to 100 words, along with their 768-dimensional embedded vectors. The retrieval process involves converting a query into a DPR embedding and finding the top k vectors with the closest L2 distances. For both systems, we retrieve the top 5 results, concatenate them with the original query, and feed them to the LLM reader.
Generation  Although different prompting strategies may influence the performance of the question-answering task, this aspect is not the primary focus of our study, so we adhere to the same answer prompts used in the RRR Ma et al. (2023) framework. The prompts we used are detailed in Table 1.
Trainable Scheme  For Trainable ERRR, we employ T5-Large Raffel et al. (2020), an open-source model with 770 million parameters, as the query optimizer. We fine-tune this student model using knowledge distillation from GPT-3.5-Turbo. The distillation dataset was assembled by selecting questions from training sets of each QA dataset, with GPT-3.5-Turbo generating the responses under identical settings utilized in the frozen scheme. We also devised a short eliciting prompt, "Rewrite better search queries to acquire or validate the knowledge needed for the question:", serving as an instruction prefix to guide T5 to adapt to the task. To ensure optimal task-specific outcomes, separate T5 models were trained with 3 epochs for each QA dataset, with a learning rate of 1e-4 and a batch size of 4.
AmbigQA PopQA HotPotQA
Methods EM F1 EM F1 EM F1
Direct 0.391 0.4996 0.392 0.4289 0.311 0.4178
@@ 4.4 Result
Frozen ERRR Trainable ERRR ReAct Self-RAG
Cost \$0.62 \$0.53 \$1.05 \$1.65
@@
The experimental results across three datasets and two retrieval tools are presented in Table 2 and Table 3. The Frozen ERRR framework consistently outperforms all baseline methods—Direct, Frozen RRR, and Trainable RRR—regardless of the retrieval system used. These results highlight the effectiveness of addressing the pre-retrieval information gap, demonstrating ERRR’s adaptability across diverse retrieval systems and datasets. Furthermore, the Trainable ERRR framework achieves even better performance, surpassing all baselines and its teacher model (GPT-3.5 Turbo) across all three datasets. We attribute this improvement to the distillation process, which enables the student model (fine-tuned T5) to generalize better by focusing on critical features while filtering out irrelevant information. This distilled representation allows the model to adapt more effectively to specific query optimization tasks, potentially compressing and refining the teacher’s insights into a more efficient form.
The impact of the ERRR framework is more pronounced in web search retrieval systems, as evidenced by the greater performance enhancement observed in Table 2 compared to dense retrievers in Table 3. This is likely due to the higher quality and broader knowledge span of web-based retrieval systems compared to the static 2018 Wikipedia corpus used for dense retrieval. Notably, the results show that both Frozen RRR and Trainable RRR underperform the Direct method in the PopQA and HotPotQA datasets when using dense retrieval. This underperformance can be attributed to the low-quality results retrieved from the outdated and limited corpus, which includes only Wikipedia passages of constrained length and scope. These limitations lead to an increased retrieval of irrelevant documents, distracting the Large Language Model (LLM) from answering questions correctly.
In contrast, ERRR demonstrates resilience under such conditions. By optimizing queries to align with the LLM’s informational needs, ERRR reduces the retrieval of irrelevant passages, mitigating distractions caused by lower-quality retrieval. This robustness is particularly valuable when operating on suboptimal document collections, as it ensures performance gains even in challenging retrieval scenarios. A detailed case study, provided in Appendix A, further illustrates how ERRR generates precise queries that enhance retrieval effectiveness and improve final answers, even when the retrieved content includes inaccuracies.
@@ 4.5 Cost and Latency
Given our method’s emphasis on a conventional single-turn pipeline, it demonstrates superior performance in terms of cost and latency when compared to certain advanced and iterative RAG frameworks. To underscore the cost-efficiency and flexibility of our approach, we conducted a comparative analysis with ReAct Yao et al. (2022) and Self-RAG Asai et al. (2023). This experiment was carried out on 200 randomly selected questions from HotPotQA. The results presented in Table 4 highlight that while still maintaining commendable performance, Frozen ERRR exhibits significantly lower costs, faster processing times, and greater efficiency than other iterative frameworks. Moreover, Trainable ERRR has the potential to further reduce costs, particularly for large datasets, by leveraging an already fine-tuned query optimizer, thereby saving on an additional LLM call to GPT-3.5-Turbo.
@@ 5 Conclusion
In this paper, we present Extract-Refine-Retrieve-Read (ERRR) framework for Retrieval-Augmented Generation (RAG) systems. The ERRR framework is designed to optimize queries, aligning them closely with the specific informational needs of large language models (LLMs) to enhance retrieval augmentation effectiveness. Our experimental results demonstrate that our method surpasses both the naive LLM and native query rewriting framework Rewrite-Retrieve-Read on benchmark datasets such as AmbigQA Min et al. (2020), PopQA Mallen et al. (2022) and HotPotQA Yang et al. (2018), utilizing both web search tools and a dense retriever with local static corpus. It demonstrated ERRR’s remarkable adaptability across a variety of settings, data sources, and retrieval systems. This flexibility ensures that ERRR can be effectively implemented in diverse operational environments, making it a potential and adaptable component for inclusion in more advanced RAG systems. Additionally, we have developed and implemented a trainable scheme for the ERRR framework. This approach is both cost-effective and efficient as it relies on only a fine-tuned T5 model trained on a moderately sized dataset and surpasses the performance of the frozen GPT-3.5-Turbo.
@@ 6 Limitation
We acknowledge that we recognize the existence of more sophisticated Retrieval-Augmented Generation (RAG) approaches such as Self-RAGAsai et al. (2023) and CRAGYan et al. (2024). These advanced systems typically require iterative invocations of the entire pipeline to refine their answers, resulting in exceptionally high computational demands. Due to computational constraints within our study, we focused solely on scenarios that operate in a single-turn manner, wherein each module is invoked only once.
Additionally, our model does not employ any reinforcement learning techniques to enhance the performance of the supervised fine-tuned model. This decision was driven by resource limitations and observed sub-optimal performance when training with a small portion of the dataset using Proximal Policy Optimization (PPO) Schulman et al. (2017), which constrained the potential upper limit of our model’s performance.